# Breadcrumbs Reasoning: Memory-Efficient Reasoning with Compression Beacons

### v1

@@ Abstract
The scalability of large language models for long-context reasoning is severely constrained by the linear growth of their Transformer key-value cache, which incurs significant memory and computational costs. We posit that as a model generates reasoning tokens, the informational value of past generated tokens diminishes, creating an opportunity for compression. In this work, we propose to periodically compress the generation KV cache with a learned, special-purpose token and evict compressed entries. We train the model to perform this compression via a modified joint distillation and reinforcement learning (RL) framework. Our training method minimizes overhead over the conventional RL process, as it leverages RL outputs for distillation. Empirically, our method achieves a superior memory–accuracy Pareto frontier compared to both the model without cache compression and training-free compression techniques.
@@ 1 Introduction
Reasoning through token generation allows large language models (LLMs) to solve arbitrarily complex problems with a fixed depth architecture (Merrill & Sabharwal, 2023), by scaling the compute invested through the generation of more tokens (i.e., test-time scaling) (Snell et al., 2024). This practice carries high computational costs, because of the self-attention design of Transformers (Bahdanau et al., 2014; Vaswani et al., 2017; Keles et al., 2023). Not only it relies on simply generating many more tokens, but later tokens require computation over the representations of all previous tokens, incurring higher time complexity and necessitating increasing memory costs.
@@
The key to our approach is to substitute previously computed key-value (KV) cached representations with significantly more compact representations, inspired by how activation beacons are used for long-context compression (Zhang et al., 2025). We train these representations to contain the information from past tokens that is necessary for continuing the reasoning process to solve the task the model is given, allowing us to evict from the KV cache most previously computed representations. The challenge is to combine the training of these beacons into the reinforcement learning (RL) (Sutton & Barto, 2018) process that makes length-based reasoning possible, a fundamentally different process than the pre-training that enables conventional long-form generation. We design a joint RL-distillation approach, where we train the original non-compression policy using the conventional RL process with a verifier for reward computation, and concurrently distill it into a policy that jointly compresses and reasons.
We evaluate our method, Breadcrumbs Reasoning (Figure 1), on the Qwen2.5-1.5B and Phi-4-Mini models across three challenging reasoning benchmarks: Countdown (Gandhi et al., 2024), LinSys, and StarGraph (Bachmann & Nagarajan, 2024). We compare against a strong, uncompressed teacher policy trained with RL, as well as two training-free cache eviction baselines, TOVA (Oren et al., 2024) and StreamingLLM (Xiao et al., 2023). Our experiments reveal several key findings. Demonstrating a clear Pareto improvement, Breadcrumbs Reasoning enables effective test-time scaling by generating longer reasoning chains to match or exceed the teacher’s accuracy within a fixed memory budget, while still retaining 65.1–89.8% of the original performance when using 2–32x less memory at a fixed generation length. In contrast, the training-free baselines consistently underperform, stressing the necessity of a learned compression scheme for complex reasoning. We also validate our training strategy, showing that our joint RL-distillation approach matches or outperforms a more complex two-stage training pipeline, confirming its efficiency. Our code is available at https://github.com/lil-lab/breadcrumbs-reasoning.
@@ 2 Related Work and Background
Generation in Transformers-based LLMs requires reasoning over and storing a key-value (KV) cache. This entails high memory (i.e., space) and time costs, which increase as the context (i.e., the number of previous tokens) increases. Therefore long-form generation costs suffer not only from the fundamental need to generate more tokens via more steps, but also from the increasing cost of each such step. KV compression is a solution avenue that is receiving significant research attention (LI et al., 2025).
@@
An important thread within this compression literature is training models to perform KV cache compression. Nawrot et al. (2024) train models to compute importance scores, which are then used to store averaged KV cache entries instead of the original entries. Other methods train the Transformer-based LLMs themselves to summarize past KV entries (Chevalier et al., 2023; Zhang et al., 2025), so more compact representations can be retained. Our approach is inspired by the activation beacons method (Zhang et al., 2025), but with significant simplifications and adaptation for reasoning. We do away with the chunk and sub-chunk distinction, and eliminate the addition of specialized attention mechanisms. Rather, we adopt a flat segmentation into blocks, use the standard Transformer attention mechanism, and remove KV cache entries every time a beacon is processed (e.g., ex-ante). These modifications do not only simplify the implementation, but also allow for immediate eviction of cache entries, instead of a delayed one. More broadly, a drawback of these learned methods is that they require fine-tuning on a considerable amount of general-purpose pre-training data. We design a joint reasoning-compression training approach, which adds minimal overhead over the existing reasoning training processes.
@@
An alternative to training-based methods are training-free methods that perform compression at generation time. They can be divided into two main categories. The first is that of sliding-window approaches, which limit the KV cache by only including a sliding window plus an additional subset of tokens. Particularly simple and effective is StreamingLLM (Xiao et al., 2023), which finds that including a few initial sink tokens in addition to the window recovers most of the uncompressed performance. The second type does not constrain window tokens to remain in the cache, but rather tries to select the empirically more important for attention computations, as in H2O (Zhang et al., 2023) or TOVA (Oren et al., 2024).
@@
An orthogonal approach to reducing the KV cache size is reducing Chain-of-Thought reasoning length (Kang et al., 2025; Ma et al., 2025; Shen et al., 2025; Yan et al., 2025; Munkhbat et al., 2025; Xia et al., 2025). These methods primarily aim to directly shorten reasoning traces. This is distinct from our objective of dynamic KV cache compression, which focuses on extracting critical information to manage cache sizes effectively during the reasoning process itself. KV cache compression methods, like ours, could be applied in combination with those methods to achieve even higher levels of efficiency.
@@ 3.1 Breadcrumbs Reasoning
Our compression scheme uses the Transformer architecture itself for both compression and reasoning. We generate tokens following the same procedure as a vanilla Transformer-based language model, but periodically compute compressed representations of past KV cache entries, and evict these entries from the cache. Algorithm 1 describes the process. We use a special token $b$ to mark when the model should compute compressed representations of past tokens, and input this token every $c$ tokens, where $c$ is the target compression ratio. The KV cache entries for the token $b$ form the compressed representation, and we drop the entries for previous $c$ tokens. Immediately after the beacon $b$, we force the next input token to be the last sampled token before $b$ was given as input. This is equal to continuing conventional generation. Figure 1 visualizes this process to illustrate the space savings. Roughly speaking, this process leaves a trace of “reasoning breadcrumbs” behind, instead of long, detailed, and eventually irrelevant information.
@@ 3.2 Joint RL-distillation Training
Typically, LLMs are trained to solve reasoning tasks through RL (Shao et al., 2024; DeepSeek-AI, 2025; Lambert et al., 2025). Applying this process as is to a breadcrumbs reasoning policy is technically possible, but is unlikely to lead to effective learning. This is because, before training, the model does not have compression ability, so incorporating the compression token and cache eviction will significantly damage its functionality, and it will observe no positive reward during RL.
@@
For parallel and efficient training, at training time $\pi_{\rm BR}$ does not execute the KV cache removal strategy (Section 3.1), but simulates it by masking compressed tokens. Figure 2 visualizes the attention mask that results from the training masking pattern.
@@ 4 Experimental Setup
  • Countdown (Gandhi et al., 2024): the task input is a tuple of numbers, and the goal is to create a sequence of arithmetic operations using a subset of the numbers to equal a target number. While this task requires the model to avoid repeating previous attempts, or it can enter an endless loop, the individual guesses are largely independent of one another.
@@
  • StarGraph (Bachmann & Nagarajan, 2024): the model is given a list of directed edges of a star graph (i.e., a graph with multiple branches of a fixed length all expanding from a central node) and a target end node. The model must output the full sequence of edges to get to the target node from the central node. This is a task auto-regressive models such as Transformers naturally struggle with, as observed by Bachmann & Nagarajan (2024) and Hu et al. (2025).
@@
We utilize two models for our experiments: Qwen2.5-1.5B-Instruct (Team, 2025) and Phi-4-mini-instruct (Microsoft, 2025). For Qwen2.5-1.5B-Instruct, we train for 1000 steps for all tasks, while for Phi-4-mini-instruct, we train for 200 steps for StarGraph and LinSys and 1000 steps for Countdown. We use a batch size of 256 for both models. We use PPO (Schulman et al., 2017) as the RL algorithm for $\pi_{RL}$. We experiment with compression ratios $c$ of 2, 4, 8, 16, and 32 for breadcrumbs reasoning. The reward for an incorrect or not present answer is 0.0, a correct format of the final answer but an incorrect value gets a 0.1 reward, and a correct response gets a 1.0 reward.
@@
We compare our approach against two primary training-free baselines applied to $\pi_{\rm RL}$: StreamingLLM (Xiao et al., 2023) and TOVA (Oren et al., 2024). Similar to our joint training setup, they also do not require more data than what $\pi_{\rm RL}$ was trained on, and they also delete entries from the KV cache. [^1] We omit comparing to methods such as QUEST (Tang et al., 2024), because they are not focused on compression, but rather smart management of the GPU memory, an orthogonal approach to compression. For a fair comparison, we adapt these methods to use an increasingly large KV cache size during generation, matching the memory footprint of our approach. This is to avoid potentially penalizing them with a smaller static cache size. In particular, given a compression ratio $c$, we allow the cache to grow by 1 every $c$ generated tokens. We set the sliding window of StreamingLLM to $c$ to match the budget given to our approach. We also set the sink size of StreamingLLM to the entire question. For TOVA, we set the initial size of the KV cache to the number of question tokens plus $c$, so that the first compression would only happen after at least $c$ tokens, similar to our method. For both methods, we test the same $c$ as for our policies. We also study the effect of two-step training, by first training the RL policy $\pi_{\rm RL}$ and only later generating data for distillation. In both cases, we use the same number of data samples and training steps (256 samples per step, for 1k steps).
@@ 5 Results
Qwen Phi Qwen Phi Qwen Phi
Method $c$ Accc AUAC Accc AUAC Accc AUAC Accc AUAC Accc AUAC Accc AUAC
Teacher - 0.598 0.377 0.633 0.433 0.918 0.276 0.898 0.142 0.902 0.513 0.848 0.498
Breadcrumbs 2 0.605 0.465 0.609 0.506 0.730 0.451 0.652 0.373 0.957 0.726 0.812 0.636
4 0.613 0.533 0.625 0.557 0.656 0.532 0.539 0.421 0.969 0.845 0.832 0.739
8 0.613 0.555 0.613 0.581 0.410 0.368 0.363 0.322 0.973 0.906 0.836 0.786
16 0.574 0.539 0.625 0.604 0.367 0.347 0.219 0.205 0.957 0.918 0.812 0.785
32 0.535 0.513 0.613 0.597 0.297 0.286 0.195 0.187 0.957 0.926 0.816 0.792
TOVA 2 0.574 0.447 0.230 0.196 0.000 0.000 0.000 0.000 0.664 0.526 0.801 0.621
4 0.289 0.272 0.066 0.064 0.000 0.000 0.000 0.000 0.457 0.417 0.562 0.504
8 0.172 0.167 0.051 0.050 0.000 0.000 0.000 0.000 0.445 0.424 0.457 0.433
16 0.188 0.183 0.047 0.046 0.000 0.000 0.000 0.000 0.430 0.414 0.453 0.436
32 0.207 0.199 0.098 0.094 0.000 0.000 0.000 0.000 0.441 0.423 0.395 0.378
StreamingLLM 2 0.012 0.007 0.016 0.014 0.000 0.000 0.000 0.000 0.055 0.038 0.180 0.137
4 0.023 0.015 0.031 0.018 0.000 0.000 0.000 0.000 0.055 0.032 0.020 0.017
8 0.027 0.024 0.109 0.060 0.000 0.000 0.000 0.000 0.031 0.023 0.031 0.029
16 0.051 0.049 0.156 0.150 0.000 0.000 0.000 0.000 0.047 0.041 0.098 0.089
32 0.094 0.090 0.312 0.300 0.000 0.000 0.000 0.000 0.117 0.105 0.102 0.097
Table 1: Model performance on long-context reasoning tasks for a fixed generation cache size. We report accuracy given a maximum cache size of 1,000 entries (Accc) and Area Under the Accuracy Curve (AUAC). $c$ denotes the compression ratio.
Figure 3: Accuracy vs. KV Cache Size Breadcrumbs Reasoning retains most of the teacher’s performance while using significantly fewer KV cache entries, even outperforming the teacher when setting a fixed maximum KV cache size in Countdown and StarGraph.
We compare the Breadcrumbs Reasoning (BR) policy to an uncompressed RL policy and to the baselines in two different configurations. We first compare results with a set maximum cache size (Table 1 and Figure 3). We also compare with a fixed maximum number of generation tokens (Table 2), the same used during training of $\pi_{\rm RL}$.
@@
Table 1 shows accuracy performance with a fixed maximum cache budget of 1,000 steps, as well as the accuracy area under the curve (AUAC) with varying the maximum cache size up to 1,000. Figure 3 shows the curves. Across most settings, BR recovers most of the performance of the teacher at a much lower memory cache budget. Except for LinSys, which remains challenging, all compression ratios outperform the teacher for most of the budgets. This is because the compressed models are able to effectively accommodate more reasoning steps within the same cache budget, allowing for more aggressive test-time compute scaling. On Countdown, both BR models outperform even the teacher at maximum KV cache budget.
We measure the trade-off between accuracy and KV cache consumption by measuring the area under the accuracy-cache size curve. This metric integrates accuracy over the full range of cache sizes, and thus captures a model’s overall robustness to a variety of cache constraints. A larger AUAC indicates that a method maintains high accuracy across a wider range of cache budgets, rather than only performing well with high memory usage. We can observe that BR outperforms even the teacher policy by 3.6–92.8% for Qwen and 16.9—196.5% for Phi across the different ratios and tasks. For high compression ratios, the number of generated tokens is much higher than during training (e.g., 32,000 tokens vs 1,000 for 32x compression ratio). This is evidence that compression generalizes to much longer sequences than those observed during training.
We visualize this effect in Figure 4. In particular, Qwen and Phi improve by up to 14.5% and 5.1% across tasks after 1,000 generation tokens. The only exception is LinSys, where more generation tokens do not seem to help (as can be seen from Figure 3 too).
@@
Countdown LinSys StarGraph Average
Method $c$ Qwen Phi Qwen Phi Qwen Phi Qwen Phi
Teacher - 0.598 0.633 0.918 0.895 0.902 0.848 0.806 0.793
Breadcrumbs 2 0.559 0.590 0.707 0.641 0.906 0.777 0.724 0.669
4 0.578 0.574 0.648 0.516 0.883 0.797 0.703 0.629
8 0.508 0.574 0.395 0.352 0.895 0.793 0.599 0.573
16 0.430 0.594 0.359 0.207 0.867 0.785 0.552 0.529
32 0.438 0.570 0.289 0.195 0.891 0.781 0.539 0.516
TOVA 2 0.535 0.219 0.000 0.000 0.648 0.773 0.395 0.331
4 0.289 0.066 0.000 0.000 0.457 0.559 0.249 0.208
8 0.172 0.051 0.000 0.000 0.441 0.453 0.204 0.168
16 0.188 0.047 0.000 0.000 0.430 0.453 0.206 0.167
32 0.207 0.098 0.000 0.000 0.441 0.395 0.216 0.164
StreamingLLM 2 0.008 0.016 0.000 0.000 0.043 0.156 0.017 0.057
4 0.012 0.012 0.000 0.000 0.016 0.016 0.009 0.009
8 0.012 0.039 0.000 0.000 0.012 0.031 0.008 0.023
16 0.051 0.137 0.000 0.000 0.012 0.078 0.021 0.072
32 0.094 0.297 0.000 0.000 0.047 0.102 0.047 0.133
Table 2: Model accuracy on long-context reasoning tasks for a fixed generation length. The metric shown is accuracy at a sequence length of L=1, [^0] L=1{,}000. $c$ denotes the compression ratio. In particular, KV cache compression methods use only $\frac{1}{c}$ of memory compared to the teacher to achieve the reported accuracy.

### v3

@@ Abstract
The scalability of large language models for long-context reasoning is severely constrained by the linear growth of their Transformer key-value cache, which incurs significant memory and computational costs. We posit that as a model generates reasoning tokens, the informational value of past generated tokens diminishes, creating an opportunity for compression. In this work, we propose to periodically compress the generation KV cache with a learned, special-purpose token and evict compressed entries. We train the model to perform this compression via a modified joint distillation and reinforcement learning (RL) framework. Our training method minimizes overhead over the conventional RL process, as it leverages RL outputs for distillation. Empirically, our method achieves a superior memory-accuracy Pareto frontier compared to both the model without cache compression and training-free compression techniques.
@@ 1 Introduction
Reasoning through token generation allows large language models (LLMs) to solve arbitrarily complex problems with a fixed depth architecture (merrill2023expressive), by scaling the compute invested through the generation of more tokens (i.e., test-time scaling) (testtime). This practice carries high computational costs, because of the self-attention design of Transformers (Bahdanau2014NeuralMT; NIPS2017_3f5ee243; duman-keles2023selfattention). Not only it relies on simply generating many more tokens, but later tokens require computation over the representations of all previous tokens, incurring higher time complexity and necessitating increasing memory costs.
@@
The key to our approach is to substitute previously computed key-value (KV) cached representations with significantly more compact representations, inspired by how activation beacons are used for long-context compression (zhang2025long_activation_beacon). These compressed representations can be placed at single or multiple pre-determined intervals, enabling deployment-time flexibility with a single model. We train these representations to contain the information from past tokens that is necessary for continuing the reasoning process to solve the task the model is given, allowing us to evict from the KV cache most previously computed representations. The challenge is to combine the training of these beacons into the reinforcement learning (RL) (10.5555/3312046) process that makes length-based reasoning possible, a fundamentally different process than the pre-training that enables conventional long-form generation. We design a joint RL-distillation approach, where we train the original non-compression policy using the conventional RL process with a verifier for reward computation, and concurrently distill it into a policy that jointly compresses and reasons.
We evaluate our method, Breadcrumbs Reasoning (Figure 1), on the Qwen2.5-1.5B and Phi-4-Mini models across three challenging reasoning benchmarks: Countdown (gandhi2024streamsearchsoslearning), LinSys, and StarGraph (bachmann2024pitfalls). We compare against a strong, uncompressed teacher policy trained with RL, as well as four training-free cache eviction baselines: PyramidKV (cai2025pyramidkv), SnapKV (NEURIPS2024_28ab4182), TOVA (oren2024transformers), and StreamingLLM (xiao2023efficient). Our experiments reveal several key findings. Demonstrating a clear Pareto improvement, Breadcrumbs Reasoning enables effective test-time scaling by generating longer reasoning chains to match or exceed the teacher’s accuracy within a fixed memory budget, while still retaining 65.1–89.8% of the original performance when using 2–32x less memory at a fixed generation length. In contrast, the training-free baselines consistently underperform, stressing the necessity of a learned compression scheme for complex reasoning. We further show that it is possible to train a single model to support multiple compression ratios, even outperforming ratio-specific models. This allows for dynamic inference-time adjustment, where the compression ratio can be chosen based on the memory budget and required performance for each user query. We also validate our training strategy, showing that our joint RL-distillation approach matches or outperforms a more complex two-stage training pipeline, confirming its efficiency. Our code is available at https://github.com/lil-lab/breadcrumbs-reasoning.
@@ 2 Related Work and Background
Generation in Transformers-based LLMs requires reasoning over and storing a key-value (KV) cache. This entails high memory (i.e., space) and time costs, which increase as the context (i.e., the number of previous tokens) increases. Therefore long-form generation costs suffer not only from the fundamental need to generate more tokens via more steps, but also from the increasing cost of each such step. KV compression is a solution avenue that is receiving significant research attention (li2025a).
@@
An important thread within this compression literature is training models to perform KV cache compression. nawrot2024dynamic train models to compute importance scores, which are then used to store averaged KV cache entries instead of the original entries. Other methods train the Transformer-based LLMs themselves to summarize past KV entries (NEURIPS2023_3d77c6dc; chevalier2023adapting; zhang2025long_activation_beacon), so more compact representations can be retained. Our approach is inspired by the activation beacons method (zhang2025long_activation_beacon), but with significant simplifications and adaptation for reasoning. We do away with the chunk and sub-chunk distinction, and eliminate the addition of specialized attention mechanisms. Rather, we adopt a flat segmentation into blocks, use the standard Transformer attention mechanism, and remove KV cache entries every time a beacon is processed (e.g., ex-ante). These modifications do not only simplify the implementation, but also allow for immediate eviction of cache entries, instead of a delayed one. More broadly, a drawback of these learned methods is that they require fine-tuning on a considerable amount of general-purpose pre-training data. We design a joint reasoning-compression training approach, which adds minimal overhead over the existing reasoning training processes.
@@
An alternative to training-based methods are training-free methods that perform compression at generation time. They can be divided into two main categories. The first is that of sliding-window approaches, which limit the KV cache by only including a sliding window plus an additional subset of tokens. Particularly simple and effective is StreamingLLM (xiao2023efficient), which finds that including a few initial sink tokens in addition to the window recovers most of the uncompressed performance. The second type does not constrain window tokens to remain in the cache, but rather tries to select the empirically more important for attention computations, as in H2O (zhang2023h2o) or TOVA (oren2024transformers). Other approaches focus on prefill compression, reducing tokens by using the attention scores of the window of most recent token of the entire prefilling prompt (cai2025pyramidkv; NEURIPS2024_28ab4182). While these latter methods are designed for prefill and do not compress generation tokens directly, they can be applicable to our scenario by applying them repeatedly to compress at generation time.
@@
An orthogonal approach to reducing the KV cache size is reducing Chain-of-Thought reasoning length (aggarwal2025l; kang2025c3ot; ma2025cot; shen2025dast; yan2025inftythink; munkhbat2025self; xia2025tokenskip). These methods primarily aim to directly shorten reasoning traces. This is distinct from our objective of dynamic KV cache compression, which focuses on extracting critical information to manage cache sizes effectively during the reasoning process itself. KV cache compression methods, like ours, could be applied in combination with those methods to achieve even higher levels of efficiency.
@@ 3.1 Breadcrumbs Reasoning
Our compression scheme uses the Transformer architecture itself for both compression and reasoning. We generate tokens following the same procedure as a vanilla Transformer-based language model, but periodically compute compressed representations of past KV cache entries, and evict these entries from the cache. Algorithm 1 describes the process. We add a special token $b$ to the model vocabulary and embedding matrix, to mark when the model should compute compressed representations of past tokens. We input this token every $c$ tokens, where $c$ is the target compression ratio. The KV cache entries for the token $b$ form the compressed representation, and we drop the entries for previous $c$ tokens. Immediately after the beacon $b$, we force the next input token to be the last sampled token before $b$ was given as input. This is equal to continuing conventional generation. Figure 1 visualizes this process to illustrate the space savings. Roughly speaking, this process leaves a trace of “reasoning breadcrumbs” behind, instead of long, detailed, and eventually irrelevant information.
@@ 3.2 Joint RL-distillation Training
Typically, LLMs are trained to solve reasoning tasks through RL (deepseek-math; deepseekai2025deepseekr1incentivizingreasoningcapability; lambert2025tulu). Applying this process as is to a breadcrumbs reasoning policy is technically possible, but is unlikely to lead to effective learning. This is because, before training, the model does not have compression ability, so incorporating the compression token and cache eviction will significantly damage its functionality, and it will observe no positive reward during RL.
@@
For parallel and efficient training, at training time $\pi_{\rm BR}$ does not execute the KV cache removal strategy (Section 3.1), but simulates it by masking compressed tokens. Figure 2 visualizes the attention mask that results from the training masking pattern. $\pi_{\rm BR}$ learns to use the beacon token $b$ and to compress by aligning its next token distribution to the next token distribution of a policy without compression, $\pi_{\rm RL}$. Thanks to the attention mask, the model at training time needs to learn how to retain useful information through the beacons’ activations, as future tokens could not otherwise use it.
@@ 4 Experimental Setup
  • Countdown (gandhi2024streamsearchsoslearning): the task input is a tuple of numbers, and the goal is to create a sequence of arithmetic operations using a subset of the numbers to equal a target number. While this task requires the model to avoid repeating previous attempts, or it can enter an endless loop, the individual guesses are largely independent of one another.
@@
  • StarGraph (bachmann2024pitfalls): the model is given a list of directed edges of a star graph (i.e., a graph with multiple branches of a fixed length all expanding from a central node) and a target end node. The model must output the full sequence of edges to get to the target node from the central node. This is a task auto-regressive models such as Transformers naturally struggle with, as observed by bachmann2024pitfalls and hu2025the.
@@
We utilize two models for our experiments: Qwen2.5-1.5B-Instruct (qwen2.5) and Phi-4-mini-instruct (microsoft2025phi4mini). In BR, we add to an additional token, the beacon $b$, to the vocabulary of these models and initialize its embeddings with the average of all other embeddings. For Qwen2.5-1.5B-Instruct, we train for 1000 steps for all tasks, while for Phi-4-mini-instruct, we train for 200 steps for StarGraph and LinSys and 1000 steps for Countdown. We use a batch size of 256 for both models. We use PPO (schulman2017proximalpolicyoptimizationalgorithms) as the RL algorithm for $\pi_{RL}$. We experiment with compression ratios $c$ of 2, 4, 8, 16, and 32 for breadcrumbs reasoning. We instantiate our training process in two different ways. In Sr Br (Single Ratio Breadcrumbs Reasoning), each model is trained on a single compression ratio. In Mr Br (Multi Ratio Breadcrumbs Reasoning), we train a single model on all compression ratios, by repeating each batch for all compression ratios before each model weights update. Mr Br models can be used at generation time with any desired ratio. The reward for an incorrect or not present answer is 0.0, a correct format of the final answer but an incorrect value gets a 0.1 reward, and a correct response gets a 1.0 reward.
@@
We compare our approach against four training-free baselines applied to $\pi_{\rm RL}$: PyramidKV (cai2025pyramidkv), SnapKV (NEURIPS2024_28ab4182), TOVA (oren2024transformers), and StreamingLLM (xiao2023efficient). Similar to our joint training setup, they also do not require more data than what $\pi_{\rm RL}$ was trained on, and they also delete entries from the KV cache. [^1] We omit comparing to methods such as QUEST (tang2024quest), because they are not focused on compression, but rather smart management of the GPU memory, an orthogonal approach to compression. For a fair comparison, we adapt these methods to use an increasingly large KV cache size during generation, matching the memory footprint of our approach. This is to avoid potentially penalizing them with a smaller static cache size. In particular, given a compression ratio $c$, we allow the cache to grow by 1 every $c$ generated tokens. We set the sliding window of StreamingLLM and the observation window of SnapKV and PyramidKV to $c$ to match the budget given to our approach. We also set the sink size of StreamingLLM to the entire question. For SnapKV, PyramidKV, and TOVA, we set the initial size of the KV cache to the number of question tokens plus $c$, so that the first compression would only happen after at least $c$ tokens, similar to our method. For all methods, we test the same $c$ as for our policies. We also study the effect of two-step training, by first training the RL policy $\pi_{\rm RL}$ and only later generating data for distillation. In both cases, we use the same number of data samples and training steps (256 samples per step, for 1k steps).
@@ 5 Results
Compr. Ratio 1 2 4 8 16 32 1 2 4 8 16 32 1 2 4 8 16 32
[…]
Teacher 0.598 - - - - - 0.918 - - - - - 0.902 - - - - -
[…]
Sr Br (Ours) - 0.605 0.613 0.613 0.574 0.535 - 0.730 0.656 0.410 0.367 0.297 - 0.957 0.969 0.973 0.957 0.957
Mr Br (Ours) - 0.613 0.617 0.629 0.605 0.582 - 0.711 0.629 0.473 0.414 0.328 - 0.965 0.961 0.980 0.961 0.965
[…]
PyramidKV - 0.445 0.215 0.117 0.078 0.012 - 0.000 0.000 0.000 0.000 0.000 - 0.605 0.504 0.516 0.477 0.539
SnapKV - 0.449 0.219 0.141 0.102 0.082 - 0.004 0.000 0.000 0.000 0.000 - 0.660 0.523 0.508 0.465 0.500
TOVA - 0.574 0.289 0.172 0.188 0.207 - 0.000 0.000 0.000 0.000 0.000 - 0.664 0.457 0.445 0.430 0.441
StreamingLLM - 0.012 0.023 0.027 0.051 0.094 - 0.000 0.000 0.000 0.000 0.000 - 0.055 0.055 0.031 0.047 0.117
Phi
Countdown LinSys StarGraph
Compr. Ratio 1 2 4 8 16 32 1 2 4 8 16 32 1 2 4 8 16 32
[…]
Teacher 0.633 - - - - - 0.898 - - - - - 0.848 - - - - -
[…]
Sr Br (Ours) - 0.609 0.625 0.613 0.625 0.613 - 0.652 0.539 0.363 0.219 0.195 - 0.812 0.832 0.836 0.812 0.816
Mr Br (Ours) - 0.586 0.594 0.613 0.586 0.625 - 0.668 0.582 0.461 0.297 0.270 - 0.809 0.805 0.812 0.758 0.766
[…]
PyramidKV - 0.484 0.238 0.184 0.020 0.000 - 0.016 0.000 0.000 0.000 0.000 - 0.637 0.402 0.305 0.336 0.367
SnapKV - 0.469 0.230 0.152 0.062 0.012 - 0.008 0.000 0.000 0.000 0.000 - 0.652 0.465 0.344 0.340 0.312
TOVA - 0.230 0.066 0.051 0.047 0.098 - 0.000 0.000 0.000 0.000 0.000 - 0.801 0.562 0.457 0.453 0.395
StreamingLLM - 0.016 0.031 0.109 0.156 0.312 - 0.000 0.000 0.000 0.000 0.000 - 0.180 0.020 0.031 0.098 0.102
(a) Model performance on long-context reasoning tasks for a fixed generation cache size. We report accuracy given a maximum cache size of 1,000 entries. Bold indicates the best result; underlined indicates the second best.
Figure 3: Accuracy vs. KV Cache Size Multi Ratio Breadcrumbs Reasoning retains most of the teacher’s performance while using significantly fewer KV cache entries, even outperforming the teacher when setting a fixed maximum KV cache size in Countdown and StarGraph.
We compare the Breadcrumbs Reasoning policy to an uncompressed RL policy and to the baselines in two different configurations. We first compare results with a set maximum cache size (Section 5, Appendix E, and Figure 6). We also compare with a fixed maximum number of generation tokens (Section 5), the same used during training of $\pi_{\rm RL}$.
@@
Section 5 shows accuracy performance with a fixed maximum cache budget of 1,000 steps. Appendix E in Appendix E provides the accuracy area under the curve (AUAC) while varying the maximum cache size up to 1,000. Figure 6 plots accuracy vs. cache size. Across most settings, BR recovers most of the performance of the teacher at a much lower memory cache budget, both in the single ratio and multi ratio settings. Except for LinSys, which remains challenging, all compression ratios outperform the teacher for most of the budgets. This is because the compressed models are able to effectively accommodate more reasoning steps within the same cache budget, allowing for more aggressive test-time compute scaling. On Countdown, both BR models outperform even the teacher at maximum KV cache budget.
We measure the trade-off between accuracy and KV cache consumption by measuring the area under the accuracy-cache size curve (Appendix E). This metric integrates accuracy over the full range of cache sizes, and thus captures a model’s overall robustness to a variety of cache constraints. A larger AUAC indicates that a method maintains high accuracy across a wider range of cache budgets, rather than only performing well with high memory usage. We can observe that BR outperforms even the teacher policy by 3.6–92.8% for Qwen and 10.9—219.7% for Phi across the different ratios and tasks. For high compression ratios, the number of generated tokens is much higher than during training (e.g., 32,000 tokens vs 1,000 for 32x compression ratio). This is evidence that compression generalizes to much longer sequences than those observed during training, except for LinSys. Figure 7 in the appendix visualizes this effect. Crucially, despite potentially generating significantly more tokens (up to $32\times$), BR maintains an inference time comparable to the teacher and on average faster than all training-free baselines (Section 5).
@@
Qwen
Countdown LinSys StarGraph
Compr. Ratio 1 2 4 8 16 32 1 2 4 8 16 32 1 2 4 8 16 32
[…]
Teacher 0.598 - - - - - 0.918 - - - - - 0.902 - - - - -
[…]
Sr Br (Ours) - 0.559 0.578 0.508 0.430 0.438 - 0.707 0.648 0.395 0.359 0.289 - 0.906 0.883 0.895 0.867 0.891
Mr Br (Ours) - 0.559 0.539 0.551 0.527 0.531 - 0.691 0.625 0.457 0.406 0.328 - 0.875 0.898 0.898 0.875 0.895
[…]
PyramidKV - 0.438 0.211 0.113 0.078 0.012 - 0.000 0.000 0.000 0.000 0.000 - 0.598 0.504 0.516 0.477 0.539
SnapKV - 0.441 0.219 0.141 0.102 0.082 - 0.004 0.000 0.000 0.000 0.000 - 0.660 0.523 0.508 0.465 0.500
TOVA - 0.535 0.289 0.172 0.188 0.207 - 0.000 0.000 0.000 0.000 0.000 - 0.648 0.457 0.441 0.430 0.441
StreamingLLM - 0.008 0.012 0.012 0.051 0.094 - 0.000 0.000 0.000 0.000 0.000 - 0.043 0.016 0.012 0.012 0.047
Phi
Countdown LinSys StarGraph
Compr. Ratio 1 2 4 8 16 32 1 2 4 8 16 32 1 2 4 8 16 32
[…]
Teacher 0.633 - - - - - 0.895 - - - - - 0.848 - - - - -
[…]
Sr Br (Ours) - 0.590 0.574 0.574 0.594 0.570 - 0.641 0.516 0.352 0.207 0.195 - 0.777 0.797 0.793 0.785 0.781
Mr Br (Ours) - 0.566 0.547 0.578 0.555 0.582 - 0.617 0.574 0.434 0.285 0.270 - 0.773 0.762 0.770 0.734 0.734
[…]
PyramidKV - 0.477 0.234 0.184 0.020 0.000 - 0.016 0.000 0.000 0.000 0.000 - 0.613 0.402 0.305 0.336 0.367
SnapKV - 0.469 0.230 0.152 0.062 0.012 - 0.008 0.000 0.000 0.000 0.000 - 0.652 0.465 0.344 0.340 0.312
TOVA - 0.219 0.066 0.051 0.047 0.098 - 0.000 0.000 0.000 0.000 0.000 - 0.773 0.559 0.453 0.453 0.395
StreamingLLM - 0.016 0.012 0.039 0.137 0.297 - 0.000 0.000 0.000 0.000 0.000 - 0.156 0.016 0.031 0.078 0.102
(a) Model accuracy on long-context reasoning tasks for a fixed generation length. The metric shown is accuracy at a sequence length of L=1, [^0] L=1{,}000. Bold indicates the best result; underlined indicates the second best.